# AI-Informed Model Analogs for Subseasonal-to-Seasonal Prediction

### v2

@@ 4 Conclusions
We demonstrate how an AI-informed model analog forecasting approach, previously only shown to be skillful on seasonal-to-decadal timescales, can also produce skillful deterministic and probabilistic subseasonal-to-seasonal predictions. We showcase this approach’s improvement over climatological, persistence, and traditional analog forecasting methods in both a perfect model framework and with reanalysis data for classification and continuous prediction tasks. For example, we find 49% and 19% increase in probabilistic skill testing on observational data for Tasks #1 and #3, respectively. While skill relative to climatology remains modest, especially for shorter timescales (e.g., Task #1 and Task #3), this is typical for S2S prediction. For example, the Seasonal-to-Multiyear Large Ensemble (SMYLE) prediction system, which similarly utilizes CESM2, finds zero or minimal skill (ACC <.3<.3) when predicting month 1-3 summer Midwestern U.S. temperatures [smyle].
@@
The AI-informed analog ensembles additionally provides improved prediction of extremes compared to traditional analog forecasting methods. Since extreme temperature events have a disproportionate impact on human health, agriculture, and energy/water management, improving their prediction is essential for mitigating the most dire consequences [annurev:/content/journals/10.1146/annurev-earth-071719-055228].
Like previous work [[, e.g.,]]Rader:2023aa,toride2024usingdeeplearningidentify,fernandez2025multiyeartodecadaltemperaturepredictionusing, we leverage a model analog approach rather than rely on observations alone to form our analog library. While we use a single model (CESM2) to form our analog library, [fernandez2025multiyeartodecadaltemperaturepredictionusing] experiment with using multiple models on multi-year-to-decadal timescales to both form the analog library and to train their neural network. Including more climate models or, specifically for S2S timescales, extended-range AI or dynamical weather forecasts [[, e.g.,]]lang2024aifsecmwfsdatadriven,Vitart:2022 may similarly boost skill on these shorter timescales.
Further, although we have shown that this method produces skillful S2S predictions on near observational data, the approach is still limited by the model’s, in this case CESM2’s, biases [[, e.g.,]]Pang:2024aa,Wei:2021aa,Woelfle:2019aa. Future work could also explore how to best incorporate near-observational data into the AI-based analog forecasting approach. For example, [fernandez2025multiyeartodecadaltemperaturepredictionusing] found that transfer learning from models to reanalysis improved analog prediction skill; a similar approach could be useful for learning masks on S2S timescales as well. Moreover, reanalysis data could also be incorporated into the analog library itself, perhaps offering even more realistic climate trajectories than those from a library composed solely of model data. These additions to the analog library and/or training set could further advance AI-based analog S2S forecasts, especially for regions and seasons where model biases are more pronounced.
@@ 5 Acknowledgments
J.B.L. acknowledges support from NOAA grants #NA22OAR4310621 and #NA19OAR4590151. The authors wish to thank Martin Fernandez for helpful insights for this research.
@@ 6 Data and Code Availability
CESM2-LE data is available at https://www.earthsystemgrid.org/dataset/ucar.cgd.cesm2le.output.html [Danabasoglu:2020aa]. ERA5 data is available at https://cds.climate.copernicus.eu/datasets [Hersbach:2020aa]. Code used to perform the analysis and generate the figures is available at https://github.com/jlandsbe/S2S.git.
[]Supplemental Information for AI-Informed Model Analogs for S2S Prediction
@@
We also assess the short-range S2S skill of the AI-based analog approach by classifying Week 3-4 Southern California ($32^{\circ}-37^{\circ}\mathrm{N},\ 116^{\circ}-121^{\circ}\mathrm{W}$) summer temperatures (Task #1). The three target classes (cold, neutral, and warm) are formed as in Task #2, where target temperatures are split into terciles, ensuring all classes are equally sized. We predict each 2-week period from the third week of June through the third week of September, using the learned weights in Figure S1. This mask exhibits weights that are distributed globally, yet unevenly, with noticeably increased weight around the western U.S. as well as in the North Pacific. The more diffuse weighting pattern, as compared to Tasks #1 and #2, likely reflects the noisier synoptic patterns present in the daily data used to train this mask.
@@
For monthly CESM2-LE data, we use the ensemble members listed in Table S2, while for daily CESM2-LE data, we use the ensemble members listed in Table S3. The first four numbers of the member names correspond to the chosen model start dates (varying initial climate conditions), while the last three numbers indicate the realization (small perturbations to initial conditions). While we use a mixture of ocean initializations, these differences have been found to not impact S2S prediction five centuries beyond their initialization date [Arcodia_2023].
@@
S4 Figures

### v3

@@ 4 Conclusions
We demonstrate how an AI-informed model-analog forecasting approach, previously only shown to be skillful on seasonal-to-decadal timescales, can also produce skillful deterministic and probabilistic subseasonal-to-seasonal predictions. While the goal of this work is not to outperform operational S2S forecasting systems, a limited comparison to operational models in Supplemental Section S4, indicates AI-based model-analogs are in line with operational model performance. Further, we showcase this approach’s improvement over climatological, persistence, and traditional model-analog forecasting methods in both a perfect model framework and with reanalysis data for classification and continuous prediction tasks. For example, we find 49% and 19% increase in probabilistic skill testing on observational data for Tasks #1 and #3, respectively. While skill relative to climatology remains modest, especially for shorter timescales (e.g., Task #1 and Task #3), this is typical for S2S prediction. For example, the Seasonal-to-Multiyear Large Ensemble (SMYLE) prediction system, which similarly utilizes CESM2, finds zero or minimal skill (ACC <.3<.3) when predicting month 1-3 summer Midwestern U.S. temperatures [36].
@@
The AI-informed model-analog ensembles additionally provides improved prediction of extremes compared to traditional analog forecasting methods. Since extreme temperature events have a disproportionate impact on human health, agriculture, and energy/water management, improving their prediction is essential for mitigating the most dire consequences [1].
Like previous work [[, e.g.,]]Rader:2023aa,toride2024usingdeeplearningidentify,fernandez2025multiyeartodecadaltemperaturepredictionusing, we leverage a model-analog approach rather than rely on observations alone to form our analog library. While we use a single model (CESM2) to form our analog library, [10] experiment with using multiple models on multi-year-to-decadal timescales to both form the analog library and to train their neural network. Including more climate models or, specifically for S2S timescales, extended-range AI or dynamical weather forecasts [[, e.g.,]]lang2024aifsecmwfsdatadriven,Vitart:2022 may similarly boost skill on these shorter timescales.
Further, although we have shown that this method produces skillful S2S predictions on near observational data, the approach is still limited by the model’s, in this case CESM2’s, biases [[, e.g.,]]Pang:2024aa,Wei:2021aa,Woelfle:2019aa. Future work could also explore how to best incorporate near-observational data into the AI-based model-analog forecasting approach. For example, [10] found that transfer learning from models to reanalysis improved analog prediction skill; a similar approach could be useful for learning masks on S2S timescales as well. Moreover, reanalysis data could also be incorporated into the analog library itself, perhaps offering even more realistic climate trajectories than those from a library composed solely of model data. These additions to the analog library and/or training set could further advance AI-based model-analog S2S forecasts, especially for regions and seasons where model biases are more pronounced.
@@ 5 Acknowledgments
J.B.L. acknowledges support from NOAA grants #NA22OAR4310621 and #NA19OAR4590151. The authors wish to thank Martin Fernandez for helpful insights for this research and Yan Wang for providing and processing NMME model data. The authors also wish to thank the editor and two anonymous reviewers for their helpful comments and suggestions.
@@ 6 Data and Code Availability
CESM2-LE data is available at https://www.earthsystemgrid.org/dataset/ucar.cgd.cesm2le.output.html [6]. ERA5 data is available at https://cds.climate.copernicus.eu/datasets [11]. Code used to perform the analysis and generate the figures is available at https://github.com/jlandsbe/S2S.git.
@@
We also assess the short-range S2S skill of the AI-based model-analog approach by classifying Week 3-4 Southern California ($32^{\circ}-37^{\circ}\mathrm{N},\ 116^{\circ}-121^{\circ}\mathrm{W}$) summer temperatures (Task #1). The three target classes (cold, neutral, and warm) are formed as in Task #2, where target temperatures are split into terciles, ensuring all classes are equally sized. We predict each 2-week period from the third week of June through the third week of September, using the learned weights in Figure S1. This mask exhibits weights that are distributed globally, yet unevenly, with noticeably increased weight around the western U.S. as well as in the North Pacific. The more diffuse weighting pattern, as compared to Tasks #1 and #2, likely reflects the noisier synoptic patterns present in the daily data used to train this mask.
@@
For monthly CESM2-LE data, we use the ensemble members listed in Table S2, while for daily CESM2-LE data, we use the ensemble members listed in Table S3. The first four numbers of the member names correspond to the chosen model start dates (varying initial climate conditions), while the last three numbers indicate the realization (small perturbations to initial conditions). While we use a mixture of ocean initializations, these differences have been found to not impact S2S prediction five centuries beyond their initialization date [2].
@@
S4 Model Baslines
While our work focuses on highlighting AI analogs utility for improving S2S model-analog forecasting and for understanding sources of predictability, here we offer a limited comparison to operational models to masked analogs Task #1 as a reference point. We compute the MAE skill of 3 models from the North American Multi-Model Ensemble (NMME): NCEP-CFSv2 (NCEP), ECCC-CanESM5 (CAN), and ECCC-GEM5.2-NEMO (GEM) [14]. We use the ensemble mean over 25 NCEP ensemble members’ predictions from 1982-2010 and 21 CAN and GEM ensemble members’ predictions from 1990-2023. We regrid all predictions to match the ERA5 grid, detrend, and convert to anomalies as described in Section 2.3. These models are reported at .5- and 1.5-month leads, both of which we include both in Table S4. We find that analog forecasting achieves lower skill than NCEP, but higher than CAN and GEM (Table S4). Moreover, for the 15% of most extreme forecasts, AI-based model-analog forecasting again achieves similar skill to the NMME models. AI-based model-analogs are outperformed by one model (CAN) but outperform the other two (Table S4). Although a limited comparison is done here, these results suggest that skill for AI-based model-analog forecasting is in line with operational S2S models.
Model Lead Total Skill Skill on Most Extreme Months
Analogs 1 Month .068 0.24
NCEP 0.5 Months 0.14 0.23
NCEP 1.5 Months 0.085 0.19
CAN 0.5 Months <0<0 0.30
CAN 1 Month <0<0 0.29
GEM 0.5 Months <0<0 0.23
GEM 1.5 Months 0.015 0.0015
Table S4: Skill scores for the AI-based analog approach as compared to operational models (where a higher score indicates a more skillful forecast relative to climatology). We find that analog forecasting (bolded) has performance in line with NMME models for Task #1. It outperforms two models and is outperformed by one, both for total skill and skill on the most extreme months.
S5 Figures